(a) An open macroeconomic model for a hypothetical economy is represented as follows

Y= C0 +Io+Go+X0-M, M=mo+m1yd,C=co+c1yd, T=tY and Yd=Y-T

Show that equal change in tax and government expenditure are expansionary to the economy
Derive the equilibrium level of savings in the economy above
Derive the investment multiplier
1
Expert's answer
2016-01-19T08:28:40-0500
Y= C0 +Io+Go+X0-M, M=mo+m1yd,C=co+c1yd, T=tY and Yd=Y-T
a) The equal change in tax and government expenditure are expansionary for the economy, because if the government expenditure increases (Go to G1), the GDP will increase too, as Y= C0 +Io+Go+X0-M. So, the equal change in tax and government expenditure will have expansionary effect.
b) In the equilibrium savings are equal to investment, so in our case the equilibrium level of savings is S = Io.
Investment multiplier is simply the multiplier effect of an injection of investment into an economy.
c) The investment multiplier in our case will be mi = 1/(1 - c) = 1/(1 - c1).